OpenStudy (anonymous):

500 juniors at Central High School took the Act last year> The scores were normally distributed with a mean of 24 and a standard deviation of 4. A)What percentage of scores are lower than 12? B)Approximately how many juniors score high than 30

OpenStudy (anonymous):

can you help me @vishweshshrimali5

OpenStudy (vishweshshrimali5):

Sorry statistics is not my strong point

OpenStudy (anonymous):

do you know someone who can help

OpenStudy (anonymous):

can you help @OOOPS

OpenStudy (vishweshshrimali5):

@ganeshie8 well he can help you or provide the names of the users.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@No.name

OpenStudy (anonymous):

anyone!!!!

ganeshie8 (ganeshie8):

familiar with calculating zscores ?

OpenStudy (anonymous):

hw do you do that

ganeshie8 (ganeshie8):

\[\large \text{Zscore = } \dfrac{\text{observation} - \text{mean}}{\text{standard deviation}}\]

ganeshie8 (ganeshie8):

`A)What percentage of scores are lower than 12?` start by finding the Zscore for the observation 12

OpenStudy (anonymous):

so zscore=500-24/12

OpenStudy (anonymous):

sorry it is 12=500-24/4

OpenStudy (anonymous):

i got 119 though

ganeshie8 (ganeshie8):

nope, you're given : mean = 24 standard deviation = 4 Zscore for observation 12 : \[\large \text{Zscore = } \dfrac{\text{12} - \text{24}}{\text{4}} = -3 \]

OpenStudy (anonymous):

oh

ganeshie8 (ganeshie8):

look up the area for a Zscore of -3 in your ztable

OpenStudy (anonymous):

0.0010

ganeshie8 (ganeshie8):

http://www.stat.ufl.edu/~athienit/Tables/Ztable.pdf

ganeshie8 (ganeshie8):

I am seeing 0.0013 ?

OpenStudy (anonymous):

sorry i did -3.1 by accident its 0.0013

ganeshie8 (ganeshie8):

Yes ! multiply it by 100 to get the percent : 0.0013 is the probability, which is same as 0.0013x100 = 0.13 %

ganeshie8 (ganeshie8):

So 0.13% of the juniors will score less than 12 in ACT.

ganeshie8 (ganeshie8):

we're done with partA! see if you can try partB

OpenStudy (anonymous):

so 30-24/4=1.5

OpenStudy (anonymous):

so 93.32%

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

!!!!!

ganeshie8 (ganeshie8):

small blunder

ganeshie8 (ganeshie8):

Ztable always gives you the probability for "<"

ganeshie8 (ganeshie8):

B)Approximately how many juniors score `high` than 30 since you want to know higher than 30, you need to subtract the area from 100% : percent students higher than 30 = 100% - 93.32% = 6.68%

ganeshie8 (ganeshie8):

whats 6.68% of 500 ?